OpenStudy (anonymous):

f(x)= (3x^3)/[(4-5x)^5] find the equation of the line tangent to the graph of f at x=2.

OpenStudy (angela210793):

eq of the tangent is: y-f(2)=f'(2)(x-2) Do u know how to solve it now?

OpenStudy (anonymous):

No, I get m, but not b

OpenStudy (anonymous):

for x=2, the value I get is .00823, unless I have the wrong derivative.

OpenStudy (angela210793):

it's very easy..try again,first find f(2) bu replacing 2 instead of x in the eq given.then find the derivative of the eq and find f'(2) and then replace the values u found

OpenStudy (angela210793):

[U/V]'=(U'V-UV')/V^2

OpenStudy (anonymous):

ok, so for my derivative I have (6x^2(5x+6))/((4-5x)^6)

OpenStudy (anonymous):

You made a mistake on the derivative, go to your other post, I wrote the derivative there.

OpenStudy (anonymous):

Also, depending on your teacher, but generally in these types of problems, it is not necessary to find decimal numbers, leave them in fractions.

OpenStudy (anonymous):

It's an online assignment, so it requires 4 decimal places. Are you sure about your derivative.

OpenStudy (anonymous):

Take the denominator for example, it is\[[(4-5x)^{5}]^{2}=(4-5x)^{10}\]

OpenStudy (anonymous):

why are you squaring the denominator?

OpenStudy (anonymous):

In the process of finding the derivative, you use a process call the quotient rule. It is usually written like this\[(u'v -uv')/v ^{2}\]

OpenStudy (anonymous):

\[u =3x ^{3}\]\[u'=9x ^{2}\]\[v =(4-5x)^{5}\]\[v'=-25(4-5x)^{4}\]

OpenStudy (anonymous):

Is this making any sense?

OpenStudy (anonymous):

somewhat.